OpenStudy (anonymous):

f(x)=1/3x for x=4 HELP

OpenStudy (anonymous):

f(4) = 1 / 3(4) = 1/12

OpenStudy (anonymous):

i think i got 1 1/3 which is equal to 4?

OpenStudy (anonymous):

\[\left| -x-2y \right| for x=-2 and y=3\]

OpenStudy (anonymous):

Wait, what's the question? Where did you get |−x−2y| for x=−2 and y=3 from?

OpenStudy (anonymous):

homework packet

OpenStudy (anonymous):

Could you post a screenshot or photo of the question? I don't quite understand what's being asked.

OpenStudy (anonymous):

evaluate \[ealuate \left| -x-2y \right| \]

OpenStudy (anonymous):

for x=-2 and y=3

OpenStudy (anonymous):

Well the '|' symbols mean you have to take the absolute value of what's inside them. So all you have to do is substitute -2 for x and 3 for y and get rid of the negative sign (if you get one). So just start by solving the equation normally, \[-x -2y\] \[=-(-2) - 2(3)\] \[= 2 - 6\] \[= -4\] But the question asks for the absolute value, so \[|-4|\] \[= 4\]